# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ A.2 $\sum_{k=0}^{\infty}k\mu(S_{R}^{k})<\infty$
Let $\mathsf{S}_{\mathrm{sde}}$ be a Borel subset of $\mathsf{S}$ such that $\mathsf{S}_{\mathrm{sde}}\subset\mathsf{S}_{\mathrm{s,i}}$. Let $\mathfrak{u}^{[1,0]}$ be the map on $\mathbb{R}^{d}\times\mathsf{S}$ such that $\mathfrak{u}^{[1,0]}(x,\mathsf{s})=\delta_{x}+\mathsf{s}$. Let $\mathbf{S}_{\mathrm{sde}}=\mathfrak{u}^{-1}(\mathsf{S}_{\mathrm{sde}})$ and $\mathsf{S}_{\mathrm{sde}}^{[1]}=(\mathfrak{u}^{[1,0]})^{-1}(\mathsf{S}_{\mathrm{sde}})$.
@@
$dX_{t}^{i}=\sigma(X_{t}^{i},\mathsf{X}_{t}^{i\diamond})dB_{t}^{i}+\mathsf{b}(X_{t}^{i},\mathsf{X}_{t}^{i\diamond})dt\ (i\in\mathbb{N}),\quad\mathbf{X}_{t}\in\mathbf{S}_{\mathrm{sde}},\forall t\in[0,\infty).$ (13.14)
Here, we set $\mathsf{X}$ and $\mathsf{X}^{i\diamond}$ as $\mathsf{X}_{t}=\sum_{i\in\mathbb{N}}\delta_{X_{t}^{i}}$ and $\mathsf{X}_{t}^{i\diamond}=\sum_{j\in\mathbb{N},\ j\not=i}\delta_{X_{t}^{j}}$. We take $\mathbf{B}=(B^{i})_{i\in\mathbb{N}}$ to be an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued Brownian motion. That is, $\{B^{i}\}_{i\in\mathbb{N}}$ are independent copies of a $d$-dimensional Brownian motion starting at the origin.
@@
belong to $\mathcal{L}^{2}$ and $\mathcal{L}^{1}$, respectively. Here, $\mathcal{L}^{p}$ is the set of all measurable $\{\mathscr{F}_{t}\}_{t\geq 0}$-adapted processes $\alpha$ such that $E[\int_{0}^{T}|\alpha(t,\omega)|^{p}dt]<\infty$ for all $T$. Here, we can and do take a predictable version of $\Phi^{i}$ and $\Psi^{i}$ (see pp 45-46 in [11]).
@@
Let $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ be a weak solution to (13.14) starting at $\mathbf{s}=\mathfrak{l}(\mathsf{s})$. Here, $\mathfrak{l}$ is the label given in (2.43). For $\mathbf{X}$, we introduce the family of finite-dimensional SDEs as follows.
Define $\sigma_{\mathsf{X}}^{m}\!:\![0,\infty)\times(\mathbb{R}^{d})^{m}\!\to\!\mathbb{R}^{d^{2}}$ and $\mathit{b}_{\mathsf{X}}^{m}\!:\![0,\infty)\times(\mathbb{R}^{d})^{m}\!\to\!\mathbb{R}^{d}$ such that, for $(u,\mathbf{v})\in(\mathbb{R}^{d})^{m}$ and $\mathsf{v}=\sum_{i=1}^{m-1}\delta_{v_{i}}\in\mathsf{S}$, where $\mathbf{v}=(v_{1},\ldots,v_{m-1})\in(\mathbb{R}^{d})^{m-1}$,
@@
Here, $\mathsf{X}_{t}^{m*}=\sum_{i=m+1}^{\infty}\delta_{X_{t}^{i}}$. For $\mathbf{s}=(s^{i})_{i\in\mathbb{N}}$, we set $\mathsf{s}_{m}^{*}=\sum_{i=m+1}^{\infty}\delta_{s^{i}}$. Recall that $\mathbf{X}_{0}=\mathbf{s}=\mathfrak{l}(\mathsf{s})$. We have $\mathsf{X}_{0}^{m*}=\mathsf{s}_{m}^{*}$ by construction. The coefficients $\sigma_{\mathsf{X}}^{m}$ and $\mathit{b}_{\mathsf{X}}^{m}$ depend on both $\mathsf{X}^{m*}$ and the label $\mathfrak{l}$. In particular, $\mathsf{X}$ gives a part of $\sigma_{\mathsf{X}}^{m}$ and $\mathit{b}_{\mathsf{X}}^{m}$. Let
$\mathbf{S}_{\mathrm{sde}}^{m}(t,\mathsf{w})=\{\mathbf{s}^{m}=(s^{1},\ldots,s^{m})\in(\mathbb{R}^{d})^{m};\mathfrak{u}(\mathbf{s}^{m})+\mathsf{w}_{t}^{m*}\in\mathsf{S}_{\mathrm{sde}}\},$
@@
$\mathbf{Y}_{0}^{m}=\mathbf{s}^{m},\quad\text{ where $\mathbf{s}^{m}=(s^{1},\ldots,s^{m})$ for $\mathbf{s}=(s^{i})\in(\mathbb{R}^{d})^{\mathbb{N}}$}.$ (13.15)
@@
We introduce the concept of a unique strong solution of (13.15). Let $\widetilde{P}^{m}=P_{\mathbf{s}}\circ(\mathbf{B}^{m},\mathbf{X}^{m*})^{-1}$ be the distribution of $(\mathbf{B}^{m},\mathbf{X}^{m*})$. Let $W^{\mathbb{N}}=C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$, $W^{m}=C([0,\infty);\mathbb{R}^{d})^{m}$, and $W_{\mathbf{0}}^{m}=\{\mathbf{w}\in W^{m};\mathbf{w}(0)=\mathbf{0}\}$. We set $\mathscr{B}_{t}(W_{\mathbf{0}}^{m}\times W^{\mathbb{N}})=\sigma[(\mathbf{v}(s),\mathbf{w}(s));0\leq s\leq t]$, $\mathscr{C}^{m}_{t}=\overline{\mathscr{B}_{t}(W_{\mathbf{0}}^{m}\times W^{\mathbb{N}})}$, and $\mathscr{C}^{m}=\overline{\mathscr{B}(W_{\mathbf{0}}^{m}\times W^{\mathbb{N}})}$, where $\overline{\cdot}$ denotes the completion with respect to $\widetilde{P}^{m}$.
@@
i $\mathbf{Y}^{m}$ is called a strong solution of (13.15) for $(\mathbf{X},\mathbf{B})$ under $P_{\mathbf{s}}$ if $(\mathbf{Y}^{m},\mathbf{B}^{m},\mathbf{X}^{m*})$ satisfies (13.15) and there exists a $\mathscr{C}^{m}$-measurable function $F_{\mathbf{s}}^{m}\!:\!W_{\mathbf{0}}^{m}\times W^{\mathbb{N}}\!\to\!W^{m}$ such that $F_{\mathbf{s}}^{m}$ is $\mathscr{C}^{m}_{t}/\mathscr{B}_{t}^{m}(W^{m})$-measurable for each $t$, and $F_{\mathbf{s}}^{m}$ satisfies $\mathbf{Y}^{m}=F_{\mathbf{s}}^{m}(\mathbf{B}^{m},\mathbf{X}^{m*})$ $P_{\mathbf{s}}$-a.s., where $\mathscr{B}_{t}^{m}(W^{m})=\sigma[\mathbf{w}(s);0\leq s\leq t]$.
@@
Definition 13.7.
For (13.15), the function $F_{\mathbf{s}}^{m}$ in Definition 13.6 i is called a strong solution starting at $\mathbf{s}^{m}$. (13.15) is said to have a unique strong solution $F_{\mathbf{s}}^{m}$ if $F_{\mathbf{s}}^{m}$ satisfies the condition in Definition 13.6 ii.
By construction, the function $F_{\mathbf{s}}^{m}$ is unique for $\widetilde{P}^{m}$-a.s.
@@
13.5. A unique strong solution of ISDEs.
In Subsection 13.5, we quote a result on unique, strong solutions of ISDEs of (13.14) from [34]. In [34], we developed a theory of SDEs to infinite-particle systems based on the classical theory in [11].
@@
We say that the uniqueness in law of weak solutions for (13.14) with initial distribution $\nu$ holds if the laws of the processes $\mathbf{X}$ and $\mathbf{X}^{\prime}$ in $W^{\mathbb{N}}$ coincide for any weak solutions $\mathbf{X}$ and $\mathbf{X}^{\prime}$ with initial distribution $\nu$.
@@
We say that the pathwise uniqueness of weak solutions of (13.14) holds if whenever $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are two weak solutions defined on the same probability space $(\Omega,\mathscr{F},P)$ with the same reference family $\{\mathscr{F}_{t}\}_{t\geq 0}$ and the same $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}$ such that $\mathbf{X}_{0}=\mathbf{X}_{0}^{\prime}$ a.s., then
@@
Definition 13.10 (pathwise uniqueness of weak solutions starting at $\mathbf{s}$).
We say that the pathwise uniqueness of weak solutions of (13.14) starting at $\mathbf{s}$ holds if the condition of Definition 13.9 holds for $\mathbf{X}_{0}=\mathbf{X}_{0}^{\prime}=\mathbf{s}$ a.s.
Let $P_{\mathrm{Br}}^{\infty}$ be the distribution of an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued Brownian motion $\mathbf{B}$ with $\mathbf{B}_{0}=\mathbf{0}$. Let $W^{\mathbb{N}}=C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$. Let $W_{\mathbf{0}}^{\mathbb{N}}=\{\mathbf{w}\in W^{\mathbb{N}}\,;\mathbf{w}_{0}=\mathbf{0}\}$. Clearly, $P_{\mathrm{Br}}^{\infty}(W_{\mathbf{0}}^{\mathbb{N}})=1$. Let $\mathscr{B}_{t}(P_{\mathrm{Br}}^{\infty})$ and $\mathscr{B}(P_{\mathrm{Br}}^{\infty})$ be the completion of $\sigma[\mathbf{w}_{s};0\leq s\leq t,\,\mathbf{w}\in W_{\mathbf{0}}^{\mathbb{N}}]$ and $\mathscr{B}(W_{\mathbf{0}}^{\mathbb{N}})$ with respect to $P_{\mathrm{Br}}^{\infty}$, respectively. Let $\mathscr{B}_{t}^{\infty}=\sigma[\mathbf{w}(s);0\leq s\leq t,\mathbf{w}\in W^{\mathbb{N}}]$.
@@
Also we call $F_{\mathbf{s}}$ itself a strong solution starting at $\mathbf{s}$.
@@
and if, for any $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}$ defined on $(\Omega,\mathscr{F},P,\{\mathscr{F}_{t}\})$ with $\mathbf{B}_{0}=\mathbf{0}$, the continuous process $F_{\mathbf{s}}(\mathbf{B})$ is a strong solution of (13.14) starting at $\mathbf{s}$. Also we call $F_{\mathbf{s}}$ a unique strong solution starting at $\mathbf{s}$.
We next present a variant of the notion of a unique strong solution.
Definition 13.13 (a unique strong solution under constraint).
@@
and if for any $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}$ defined on $(\Omega,\mathscr{F},P,\{\mathscr{F}_{t}\})$ with $\mathbf{B}_{0}=\mathbf{0}$ the continuous process $F_{\mathbf{s}}(\mathbf{B})$ is a solution of (13.14) starting at $\mathbf{s}$ satisfying $(\bullet$)). Also we call $F_{\mathbf{s}}$ a unique strong solution starting at $\mathbf{s}$ under constraint $(\bullet$)).
Let $\mathsf{S}_{\mathrm{s,i}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\}$. Let $W(\mathsf{S}_{\mathrm{s,i}})$ be the set of all $\mathsf{S}_{\mathrm{s,i}}$-valued continuous paths defined on $0,\infty)$. Let
@@
$I_{R,T}(\mathbf{w})=\sup\{i\in\mathbb{N};\min_{t\in[0,T]}\lvert w^{i}(t)\rvert\leq R\},\quad\text{where $\mathbf{w}\in W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})$}.$
Let $\mathsf{X}$ be the unlabeled path of $\mathbf{X}=(X^{i})$ given by $\mathsf{X}_{t}=\sum_{i}\delta_{X_{t}^{i}}$, where $X^{i}=\{X_{t}^{i}\}$. Recall the condition ((IFC)) introduced in Subsection 13.4. We make the assumptions.
((SIN)) $P(\mathsf{X}\in W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}}))=1$.
@@
((AC))μ $P\circ\mathsf{X}_{t}^{-1}$ is absolutely continuous with respect to $\mu$ for all $0<t<\infty$.
@@
((MF)) $P(F_{\mathbf{s}}(\mathbf{B})\in A)$ is $\overline{\mathscr{B}((\mathbb{R}^{d})^{\mathbb{N}})}^{P\circ\mathbf{X}_{0}^{-1}}$-measurable in $\mathbf{s}$ for any $A\in\mathscr{B}(W^{\mathbb{N}})$.
@@
Assume that $\mu$ is tail trivial. Assume that (13.14) has a weak solution $(\mathbf{X},\mathbf{B})$ under $P$ satisfying ((SIN)), ((NBJ)), ((AC))μ, and ((IFC)). Then, (13.14) has a family of unique strong solutions $\{F_{\mathbf{s}}\}$ starting at $\mathbf{s}$ for $P\circ\mathbf{X}_{0}^{-1}$-a.s. $\mathbf{s}$ under constraints of ((SIN)), ((NBJ)), ((AC))μ, ((IFC)), and ((MF)).
@@
Under the same assumptions as Theorem 13.4, the following hold.
i The uniqueness in law of weak solutions of (13.14) with initial distribution $\nu$ holds under constraints of ((SIN)), ((NBJ)), ((AC))μ, and ((IFC)).
ii The pathwise uniqueness of weak solutions of (13.14) with initial distribution $\nu$ holds under constraints of ((SIN)), ((NBJ)), ((AC))μ, and ((IFC)).
13.6. Cut-off functions $\chi_{\mathsf{m}}$.
We will localize the coefficients of (13.15) to prove the IFC condition. To do this, we introduce a sequence of subsets in $(\mathbb{R}^{d})^{m}\times\mathsf{S}$.
$\sigma$-compact sets $\bigcup_{q=1}^{\infty}\mathsf{K}_{Q}[\mathit{a}_{q}]$ in $\mathsf{S}$:
Let $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ be a sequence of increasing sequences $a_{\mathsf{q}}=\{a_{\mathsf{q}}(R)\}_{R\in\mathbb{N}}$ of natural numbers such that $a_{\mathsf{q}}(R)<a_{\mathsf{q}}(R+1)$ and $a_{\mathsf{q}}(R)<a_{\mathsf{q}+1}(R)$ for all $\mathsf{q},R\in\mathbb{N}$. Let
$\mathsf{K}_{Q}[\mathit{a}_{q}]$ $=\bigcap_{R=1}^{Q}\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(S_{R})\leq a_{\mathsf{q}}(R)\}\quad\text{for $Q\in\mathbb{N}\cup\{\infty\}$}.$ (13.16)
Then, $\mathsf{K}[a_{\mathsf{q}}]\subset\mathsf{K}[a_{\mathsf{q}+1}]$ for all $\mathsf{q}\in\mathbb{N}$. Note that $\mathsf{K}[a_{\mathsf{q}}]$ is a compact set in $\mathsf{S}$ for each $\mathsf{q}\in\mathbb{N}$. Let $\{a_{\mathsf{q}}^{+}(R)\}_{R\in\mathbb{N}}$ be such that $a_{\mathsf{q}}^{+}(R)=1+a_{\mathsf{q}}(R+1)$. We assume
$\mathsf{K}_{Q}[\mathit{a}_{q}]\subset\mathsf{K}_{Q}[a_{\mathsf{q}}^{+}]\subset\mathsf{K}_{Q}[a_{\mathsf{q}+1}]\quad\text{ for all $Q\in\mathbb{N}\cup\{\infty\}$}.$ (13.17)
Remark 13.1.
Note that $\mathsf{K}_{Q}[\mathit{a}_{q}]\subset\mathsf{K}_{Q}[a_{\mathsf{q}}^{+}]$ is clear because $a_{\mathsf{q}}<a_{\mathsf{q}}^{+}$. Let $C(q)$ be an increasing function such that $C(q)\to\infty$. Suppose $a_{\mathsf{q}}(R)=C(q)R^{\alpha}$ for some $\alpha>0$. Then, taking a new sequence from $\mathbf{a}=\{a_{\mathsf{q}}\}_{q\in\mathbb{N}}$ more intermittently, we can easily retake $\mathbf{a}$ such that $a_{\mathsf{q}}^{+}(R)<a_{\mathsf{q}+1}(R)$ for all $R\in\mathbb{N}$. For such $\mathbf{a}$ we obtain $\mathsf{K}_{Q}[a_{\mathsf{q}}^{+}]\subset\mathsf{K}_{Q}[a_{\mathsf{q}+1}]$. Thus, (13.17) holds.
Cut-off compact sets $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$: We introduce the approximation of $(\mathbb{R}^{d})^{m}\times\mathsf{S}$ consisting of compact sets $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$. Recall that $\mathsf{S}_{\mathrm{s,i}}$ is the set consisting of infinite configurations with no multiple points. Let $\mathbf{x}=({x}^{1},\ldots,{x}^{m})\in(\mathbb{R}^{d})^{m}$, $\mathfrak{u}(\mathbf{x})=\sum_{i=1}^{m}\delta_{{x}^{i}}$, and $\mathsf{s}=\sum_{i}\delta_{s^{i}}$. We set

### v4

@@ 13.5. Weak solution and the IFC condition to ISDEs
Let $\mathsf{S}_{\mathrm{s,i}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\}$. Let $\mathsf{S}_{\mathrm{sde}}$ be a Borel set such that $\mathsf{S}_{\mathrm{sde}}\subset\mathsf{S}_{\mathrm{s,i}}$. Let $\mathfrak{u}^{[1,0]}\!:\!\mathbb{R}^{d}\times\mathsf{S}\!\to\!\mathsf{S}$ such that $\mathfrak{u}^{[1,0]}(x,\mathsf{s})=\delta_{x}+\mathsf{s}$. Let $\mathbf{S}_{\mathrm{sde}}=\mathfrak{u}^{-1}(\mathsf{S}_{\mathrm{sde}})$ and $\mathsf{S}_{\mathrm{sde}}^{[1]}=(\mathfrak{u}^{[1,0]})^{-1}(\mathsf{S}_{\mathrm{sde}})$.
@@
$dX_{t}^{i}=\sigma(X_{t}^{i},\mathsf{X}_{t}^{i\diamond})dB_{t}^{i}+\mathsf{b}(X_{t}^{i},\mathsf{X}_{t}^{i\diamond})dt\ (i\in\mathbb{N}),\quad\mathbf{X}_{t}\in\mathbf{S}_{\mathrm{sde}}.$ (13.7)
Here, $\mathsf{X}_{t}^{i\diamond}=\sum_{j\not=i}\delta_{X_{t}^{j}}$ and $\mathbf{B}=(B^{i})_{i\in\mathbb{N}}$ is an $(\mathbb{R}^{d})^{\mathbb{N}}$-Brownian motion.
Let $\mathcal{L}^{p}$ be the set of all measurable $\{\mathscr{F}_{t}\}_{t\geq 0}$-adapted processes $\alpha$ such that $E[\int_{0}^{T}|\alpha(t,\omega)|^{p}dt]<\infty$ for all $T$.
@@
belong to $\mathcal{L}^{2}$ and $\mathcal{L}^{1}$, respectively. Here, we can and do take a predictable version of $\Phi^{i}$ and $\Psi^{i}$ (see pp 45-46 in [10]).
@@
Let $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ be a weak solution to (13.7) starting at $\mathbf{s}=\mathfrak{l}(\mathsf{s})$. Here, $\mathfrak{l}$ is the label given in (2.22). Let $\mathsf{X}$ such that $\mathsf{X}_{t}=\sum_{i}\delta_{X_{t}^{i}}$.
Define $\sigma_{\mathsf{X}}^{m}\!:\![0,\infty)\times(\mathbb{R}^{d})^{m}\!\to\!\mathbb{R}^{d^{2}}$ and $\mathit{b}_{\mathsf{X}}^{m}\!:\![0,\infty)\times(\mathbb{R}^{d})^{m}\!\to\!\mathbb{R}^{d}$ such that, for $(u,\mathbf{v})\in(\mathbb{R}^{d})^{m}$ and $\mathsf{v}=\sum_{i=1}^{m-1}\delta_{v_{i}}\in\mathsf{S}$, where $\mathbf{v}=(v_{i})_{i=1}^{m-1}\in(\mathbb{R}^{d})^{m-1}$,
@@
Here, $\mathsf{X}_{t}^{m*}=\sum_{i=m+1}^{\infty}\delta_{X_{t}^{i}}$. For $\mathbf{s}=(s^{i})_{i\in\mathbb{N}}$, we set $\mathsf{s}_{m}^{*}=\sum_{i=m+1}^{\infty}\delta_{s^{i}}$. Recall that $\mathbf{X}_{0}=\mathbf{s}=\mathfrak{l}(\mathsf{s})$. We have $\mathsf{X}_{0}^{m*}=\mathsf{s}_{m}^{*}$ by construction.
The coefficients $\sigma_{\mathsf{X}}^{m}$ and $\mathit{b}_{\mathsf{X}}^{m}$ depend on both $\mathsf{X}^{m*}$ and the label $\mathfrak{l}$. In particular, $\mathsf{X}$ gives a part of $\sigma_{\mathsf{X}}^{m}$ and $\mathit{b}_{\mathsf{X}}^{m}$. Let
$\mathbf{S}_{\mathrm{sde}}^{m}(t,\mathsf{w})=\{\mathbf{s}^{m}=(s^{i})_{i=1}^{m}\in(\mathbb{R}^{d})^{m};\mathfrak{u}(\mathbf{s}^{m})+\mathsf{w}_{t}^{m*}\in\mathsf{S}_{\mathrm{sde}}\},$
@@
$\mathbf{Y}_{0}^{m}=\mathbf{s}^{m},\quad\text{ where $\mathbf{s}^{m}=(s^{i})_{i=1}^{m}$ for $\mathbf{s}=(s^{i})\in(\mathbb{R}^{d})^{\mathbb{N}}$}.$ (13.8)
@@
Let $W=W(\mathbb{R}^{d})$ and $W_{\mathbf{0}}^{m}=\{\mathbf{w}\in W^{m};\mathbf{w}(0)=\mathbf{0}\}$.
Let $\mathscr{B}_{t}^{m}=\sigma[\mathbf{w}(s);0\leq s\leq t,\mathbf{w}\in W^{m}]$.
Let $\mathscr{B}_{t}(W_{\mathbf{0}}^{m}\times W^{\mathbb{N}})=\sigma[(\mathbf{v}(s),\mathbf{w}(s));0\leq s\leq t]$.
We set $\mathscr{C}^{m}_{t}=\overline{\mathscr{B}_{t}(W_{\mathbf{0}}^{m}\times W^{\mathbb{N}})}$, and $\mathscr{C}^{m}=\overline{\mathscr{B}(W_{\mathbf{0}}^{m}\times W^{\mathbb{N}})}$, where $\overline{\cdot}$ denotes the completion with respect to $\widetilde{P}_{\mathbf{s}}^{m}=P_{\mathbf{s}}\circ(\mathbf{B}^{m},\mathbf{X}^{m*})^{-1}$.
@@
i $\mathbf{Y}^{m}$ is called a strong solution of (13.8) for $(\mathbf{X},\mathbf{B})$ under $P_{\mathbf{s}}$ if $(\mathbf{Y}^{m},\mathbf{B}^{m},\mathbf{X}^{m*})$ satisfies (13.8) and there exists a $\mathscr{C}^{m}$-measurable function $F_{\mathbf{s}}^{m}\!:\!W_{\mathbf{0}}^{m}\times W^{\mathbb{N}}\!\to\!W^{m}$ such that $F_{\mathbf{s}}^{m}$ is $\mathscr{C}^{m}_{t}/\mathscr{B}_{t}^{m}$-measurable for each $t$, and $F_{\mathbf{s}}^{m}$ satisfies $\mathbf{Y}^{m}=F_{\mathbf{s}}^{m}(\mathbf{B}^{m},\mathbf{X}^{m*})$ $P_{\mathbf{s}}$-a.s..
@@
By construction, the function $F_{\mathbf{s}}^{m}$ is unique for $\widetilde{P}_{\mathbf{s}}^{m}$-a.s.
@@
13.6. A unique strong solution of ISDEs
@@
Uniqueness in law of weak solutions for (13.7) with initial distribution $\nu$ holds if the laws of the processes $\mathbf{X}$ and $\mathbf{X}^{\prime}$ in $W^{\mathbb{N}}$ coincide for any weak solutions $\mathbf{X}$ and $\mathbf{X}^{\prime}$ with initial distribution $\nu$.
@@
Pathwise uniqueness of weak solutions of (13.7) holds if whenever $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are two weak solutions defined on the same probability space $(\Omega,\mathscr{F},P)$ with the same reference family $\{\mathscr{F}_{t}\}_{t\geq 0}$ and the same $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}$ such that $\mathbf{X}_{0}=\mathbf{X}_{0}^{\prime}$ a.s., then
@@
Let $P_{\mathrm{Br}}^{\infty}$ be the distribution of an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued Brownian motion $\mathbf{B}$ with $\mathbf{B}_{0}=\mathbf{0}$. Let $\mathscr{B}_{t}(P_{\mathrm{Br}}^{\infty})$ and $\mathscr{B}(P_{\mathrm{Br}}^{\infty})$ be the completion of $\sigma[\mathbf{w}_{s};0\leq s\leq t,\,\mathbf{w}\in W_{\mathbf{0}}^{\mathbb{N}}]$ and $\mathscr{B}(W_{\mathbf{0}}^{\mathbb{N}})$ with respect to $P_{\mathrm{Br}}^{\infty}$, respectively. Let $\mathscr{B}_{t}^{\infty}=\sigma[\mathbf{w}(s);0\leq s\leq t,\mathbf{w}\in W^{\mathbb{N}}]$.
@@
The ISDE (13.7) admits a unique strong solution starting at $\mathbf{s}$ if there exists a $\mathscr{B}(P_{\mathrm{Br}}^{\infty})/\mathscr{B}(W^{\mathbb{N}})$-measurable mapping $F_{\mathbf{s}}\!:\!W_{\mathbf{0}}^{\mathbb{N}}\!\to\!W^{\mathbb{N}}$ such that the following hold:
@@
ii For any $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}$ defined on $(\Omega,\mathscr{F},P,\{\mathscr{F}_{t}\})$ with $\mathbf{B}_{0}=\mathbf{0}$, the continuous process $F_{\mathbf{s}}(\mathbf{B})$ is a strong solution of (13.7) starting at $\mathbf{s}$.
Definition 13.13 (a unique strong solution under constraints).
Let $(\bullet$)) be a condition. The ISDE (13.7) admits a unique strong solution starting at $\mathbf{s}$ under the constraints $(\bullet$)) if there exists a $\mathscr{B}(P_{\mathrm{Br}}^{\infty})/\mathscr{B}(W^{\mathbb{N}})$-measurable mapping $F_{\mathbf{s}}\!:\!W_{\mathbf{0}}^{\mathbb{N}}\!\to\!W^{\mathbb{N}}$ such that the following hold:
@@
ii For any $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}$ defined on $(\Omega,\mathscr{F},P,\{\mathscr{F}_{t}\})$ with $\mathbf{B}_{0}=\mathbf{0}$, the continuous process $F_{\mathbf{s}}(\mathbf{B})$ is a strong solution of (13.7) starting at $\mathbf{s}$ and satisfying $(\bullet$)).
Let $W(\mathsf{S}_{\mathrm{s,i}})$ be as in Lemma 13.1. Let
@@
$I_{R,T}(\mathbf{w})=\sup\{i\in\mathbb{N};\min_{t\in[0,T]}\lvert w^{i}(t)\rvert\leq R\},\quad\text{where $\mathbf{w}=(w^{i})\in W^{\mathbb{N}}$}.$
Let $\mathsf{X}$ be the unlabeled path of $\mathbf{X}=(X^{i})$ given by $\mathsf{X}_{t}=\sum_{i}\delta_{X_{t}^{i}}$.
[…]
@@
((AC))ν $P\circ\mathsf{X}_{t}^{-1}\ll\nu$ for all $0<t<\infty$.
Recall the condition ((IFC)) introduced in Subsection 13.5.
@@
((MF)) $P(F_{\mathbf{s}}(\mathbf{B})\in A)$ is $\overline{\mathscr{B}((\mathbb{R}^{d})^{\mathbb{N}})}$-measurable in $\mathbf{s}$ for any $A\in\mathscr{B}(W^{\mathbb{N}})$, where the completion is taken with respect to $P\circ\mathbf{X}_{0}^{-1}$}.
@@
Assume that $\nu$ is tail trivial. Assume that (13.7) has a weak solution $(\mathbf{X},\mathbf{B})$ under $P$ satisfying ((SIN)), ((NBJ)), ((AC))ν, and ((IFC)). It holds that (13.7) has a family of unique strong solutions $\{F_{\mathbf{s}}\}$ starting at $\mathbf{s}$ for $P\circ\mathbf{X}_{0}^{-1}$-a.s. $\mathbf{s}$ under the constraints ((SIN)), ((NBJ)), ((AC))ν, ((IFC)), and ((MF)).
@@
Under the same assumptions as Theorem 13.3, the following hold:
i The uniqueness in law of weak solutions of (13.7) with initial distribution $\nu$ holds under the constraints ((SIN)), ((NBJ)), ((AC))ν, and ((IFC)).
ii The pathwise uniqueness of weak solutions of (13.7) with initial distribution $\nu$ holds under the constraints ((SIN)), ((NBJ)), ((AC))ν, and ((IFC)).
13.7. Cut-off compact sets and a sufficient condition for ((IFC))
We localize the coefficients of (13.8) to prove the IFC condition. To do this, we introduce a sequence of compact subsets $\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}$ in $(\mathbb{R}^{d})^{m}\times\mathsf{S}$. Let $\mathsf{S}_{\mathrm{s,i}}$ be the set consisting of infinite configurations with no multiple points.
Let $\mathbf{x}=({x}^{1},\ldots,{x}^{m})\in(\mathbb{R}^{d})^{m}$, $\mathfrak{u}(\mathbf{x})=\sum_{i=1}^{m}\delta_{{x}^{i}}$, and $\mathsf{s}=\sum_{i}\delta_{s^{i}}$. We set